Tom is laying new sod in his yard .His front yard is 20 yd by 15 yd,and his backyard is 20 yd by 20 yd. Sod is sold by the square foot.How many square feet of sod does Tom need?: Tom is laying new sod in his yard .His front yard is 20 yd by 15 yd,and his backyard is 20 yd by 20 yd. Sod is sold by the square foot.How many square feet of sod does Tom need?
Answer by Alan3354(1937)

Frank's house needs new carpets.
the living room is 12 feet long and 13 feet wide. The dining room is 15 feet long and 11 feet wide .How many square feet of carpet will be need?
------------------
For the LR: 12*13 = 156 sf
For the DR: 15*11 = 165 sf
156+165 = 321 sq feet.
Carpet is sold and priced by the sq yard, not the square feet.